# Barriers that Programming Instructors Face While Performing Emergency Pedagogical Design to Shape Student-AI Interactions with Generative AI Tools

### v1

@@ 5.1. Open Questions for HCI Research
We synthesize our findings into open questions for future HCI research to support instructors in emergency pedagogical design.
  • Defining traits of effective student-AI interactions: Our interviewees expressed a desire to improve how students interact with GenAI tools, and often had clear ideas about undesirable behaviors (such as copy-pasting code), but struggled to articulate what constitutes a more effective or desirable student-AI interaction. How can we define a taxonomy of student-AI interactions? How can HCI research help instructors translate learning outcomes into specific, desirable and undesirable types of student-AI interactions?
  • Understanding effects of policy crosswinds: Students frequently encounter divergent policies regarding GenAI usage across different courses. How do students respond to inconsistent policies, such as one course requiring AI suggestions to be disabled while another encourages their use? What factors influence whether students adhere to these varying policies?
@@
  • Understanding student-AI behaviors with bespoke tools: Instructors who develop custom wrappers around GenAI tools theoretically have greater insight into student interactions, but often find it difficult to analyze large volumes of interaction data. How can tools help instructors make sense of student-AI interactions that are directly observable (e.g., through concept induction (lamConcept2024, 35))? How can practices from machine learning monitoring be adapted to educational contexts without requiring instructors to manage extensive infrastructure or acquire new technical expertise?
@@ 5.2. Recommendations for Academic Institutions and Funders
Figure 6. Responses to the question “In your opinion, what kinds of support would be most helpful to address faculty resistance to teaching GenAI?” Survey respondents were allowed to choose up to three options from a pre-defined list. Eight most common responses shown, with counts displayed to the right of each bar.:
When we began recruiting interviewees for this study, we initially expected that it would be straightforward to find instructors who had updated their course materials in response to GenAI, given that more than two years had passed since the release of prominent tools like ChatGPT and GitHub Copilot. However, we quickly discovered that our assumption was misplaced: it was far easier to find instructors who had only revised their course policies, rather than those who had made substantive updates to assignments, assessments, or instructional content. One possibility is that while the conversation around GenAI in education is widespread, the actual integration of GenAI into course materials remains relatively rare.
@@
Not all of the issues surfaced by our study can be addressed solely through HCI research or technical interventions; barriers like fragmented buy-in, policy crosswinds, and lack of resources are also structural or institutional in nature. In light of these findings, we offer several recommendations for academic institutions and funders to more effectively support instructors as they design better student-AI interactions. Academic institutions and departments should acknowledge that emergency pedagogical design represents additional work that instructors are taking on while still needing to manage their routine teaching responsibilities. We suggest reframing policy changes such as blanket bans on student GenAI usage as a form of emergency pedagogical design: instructors who impose such bans are still trying to reduce undesired student-AI interactions via course policy and face similar challenges in assessing whether these policies have actually succeeded in influencing student behavior. In this reframing, nearly all instructors are engaged in emergency pedagogical design of some kind, and we recommend that academic institutions reward instructors who wish to make improvements for faculty across their departments and schools. Departments should make establishing shared policies a priority to reduce the potential harms of policy crosswinds, especially on students. Institutions should provide resources for instructors who are leaders within their institutions and communities, potentially in the form of teaching reliefs, the ability to hire additional course staff, and stipends for hosting training workshops for other faculty. Where less resources are available, institutions can recognize even individual, small-scale interventions and facilitate faculty share-outs to mitigate fragmented buy-in.

### v2

@@ 5.1. Open Questions for HCI Research
As outlined by the five barriers of our research findings (Section 4), building effective courses is a challenging design problem made even more complicated as instructors are now trying to influence student-AI interactions. Some barriers, especially Implementation Challenges (Section 4.4) and Assessment Misfit (Section 4.5), suggest open questions for future HCI research to support instructors in emergency pedagogical design which we pose here for the research community.
  • Defining traits of effective student-AI interactions: Our interviewees expressed a desire to improve how students interact with GenAI tools, and often had clear ideas about undesirable behaviors (such as copy-pasting code), but struggled to articulate what constitutes more effective or desirable student-AI interaction. How can we define a taxonomy of student-AI interactions? How can HCI research help instructors translate learning outcomes into specific, desirable, and undesirable types of student-AI interactions?
  • Understanding effects of Policy Crosswinds (Section 4.3): Students frequently encounter divergent policies regarding GenAI usage across different courses. How do students respond to inconsistent policies, such as one course requiring AI suggestions to be disabled while another encourages their use? What factors influence whether students adhere to these varying policies?
@@
  • Understanding student-AI behaviors with bespoke tools: Instructors who develop custom wrappers around GenAI tools theoretically have greater insight into student interactions, but often find it difficult to analyze large volumes of interaction data. How can tools help instructors make sense of student-AI interactions that are directly observable (e.g., through concept induction (Lam et al., 2024))? How can practices from machine learning monitoring be adapted to educational contexts without requiring instructors to manage extensive infrastructure or acquire new technical expertise?
@@ 5.2. Recommendations for Academic Institutions and Funders
Figure 5. Responses to the question “In your opinion, what kinds of support would be most helpful to address faculty resistance to teaching GenAI?” Survey respondents were allowed to choose up to three options from a pre-defined list. Eight most common responses shown, with counts displayed to the right of each bar.
Bar chart showing responses to the question “In your opinion, what kinds of support would be most helpful to address faculty resistance to teaching GenAI?”. The three most common responses were: Faculty Training & Support (47 mentions); Evidence of GenAI Impact (44 mentions); and Funding & Compensation (34 mentions).
When we began recruiting interviewees for this study, we initially expected that it would be straightforward to find computing instructors who had updated their course materials in response to GenAI, given that more than two years had passed since the release of prominent tools like ChatGPT and GitHub Copilot. However, we quickly discovered that our assumption was misplaced: it was far easier to find instructors who had only revised their course policies, rather than those who had made substantive updates to assignments, assessments, or instructional content. One possibility is that while the conversation around GenAI in education is widespread, the actual integration of GenAI into course materials remains relatively rare.
@@
Not all of the issues surfaced by our study can be addressed solely through HCI research or technical interventions; barriers like Fragmented Buy-in (Section 4.2), Policy Crosswinds (Section 4.3), and Lack of Resources (Section 4.6) are also structural or institutional in nature. In light of these findings, we offer several recommendations for academic institutions and funders to more effectively support instructors as they design better student-AI interactions. Academic institutions and departments should acknowledge that emergency pedagogical design represents additional work that instructors are taking on while still needing to manage their routine teaching responsibilities. We suggest reframing policy changes such as blanket bans on student GenAI usage as a form of emergency pedagogical design: instructors who impose such bans are still trying to reduce undesired student-AI interactions via course policy and face similar challenges in assessing whether these policies have actually succeeded in influencing student behavior. In this reframing, nearly all instructors are engaged in emergency pedagogical design of some kind, and we recommend that academic institutions reward instructors who wish to make improvements for faculty across their departments and schools. Departments should make establishing shared policies a priority to reduce the potential harms of policy crosswinds, especially on students. Institutions should provide resources for instructors who are leaders within their institutions and communities, potentially in the form of teaching reliefs, the ability to hire additional course staff, and stipends for hosting training workshops for other faculty. Where less resources are available, institutions can recognize even individual, small-scale interventions and facilitate faculty share-outs to mitigate fragmented buy-in.
@@
5.3. Learning from Emergency Remote Teaching
Less than a decade ago, instructors also had to make sudden sweeping changes to their teaching in response to COVID-19. What are some lessons learned from emergency remote teaching that might help address the challenges that computing instructors now face when conducting emergency pedagogical design? In our view, one of the more helpful perspectives taken during COVID-19 was drawing a clear distinction between online learning, where a course is specifically designed for an online modality, and emergency remote teaching, where a course is offered online “in a hurry, with bare minimum resources, and scant time” (Hodges et al., 2020). In this paper, we argue for a similar distinction. While decades of research have examined how to integrate AI into pedagogy (for example, via intelligent tutoring systems (Anderson et al., 1995)), this research has also shown that effectively using AI often takes forethought and careful design. As was the case for emergency remote teaching, we view the lack of resources and time that instructors now face in emergency pedagogical design (Section 4.6) as an indicator that emergency pedagogical design should be treated as distinct from designing AI-integrated courses.
Another outcome of emergency remote teaching was that instructors became more familiar with techniques for online learning and blended learning (Singh et al., 2021) and found them valuable even after a return to in-person instruction (Adedoyin and Soykan, 2023). If this pattern also holds for emergency pedagogical design, we might expect that Fragmented Buy-in (Section 4.2) and Policy Crosswinds (Section 4.3) will become less influential as more instructors will learn about and use GenAI tools for instruction. However, this will be made more possible as we solve current problems in implementation (Section 4.4) and assessment (Section 4.5) of student-AI interactions, and thus we believe that HCI research has a critical role to play in discovering effective design practices for the future.
@@ 6. Conclusion
Acknowledgements.
We thank Google for their generous support for this project. Large language models were used to revise the text for clarity, spelling, and grammar.